OpenStudy (barinly):

HELP MEDAL OR FAN >> NEWW The position of a submarine relative to the water surface was −32 1/4 feet. A navigational maneuver decreased its depth by 15 1/2 feet. What is the position of the submarine now relative to the water surface? Enter your answer in simplest form in the box. Question 2 A fence post is 1.9 meters long. It is placed in a hole that is 0.65 meter deep. How far above the ground is the top of the fence post? Enter your answer as a decimal in the box.

OpenStudy (barinly):

@Festinger

OpenStudy (barinly):

@Festinger

OpenStudy (festinger):

What are your answers @Barinly?

OpenStudy (barinly):

Idk i think in question 2 its 1.25?

OpenStudy (barinly):

@Festinger am i right??????????????????????????? i dont know how to solve question 1

OpenStudy (festinger):

Yes, you are right for question 2. \[1.9-0.65=1.25\] For question 1 you do the same thing, but you should convert it into improper fraction for easier subtraction: \[-32\frac{1}{4}-15\frac{1}{2}=-\frac{129}{4}-\frac{31}{2}=??\]

OpenStudy (barinly):

do the denominator, have to be the same @Festinger

OpenStudy (barinly):

98/2 -- answer 49?

OpenStudy (festinger):

Yes, they have to be the same if you want to do addition or subtraction of fractions

OpenStudy (barinly):

ok

OpenStudy (barinly):

67/4 = 16/4 Right

OpenStudy (barinly):

ANSWER 16 3/4 Right @Festinger

OpenStudy (barinly):

@Festinger

OpenStudy (festinger):

No...\[-32\frac{1}{4}-15\frac{1}{2}=-\frac{129}{4}-\frac{31}{2}=-\frac{129}{4}-\frac{62}{4}=-\frac{191}{4}\] you probably forgot a minus sign.

OpenStudy (barinly):

......

OpenStudy (barinly):

@Festinger can u give me the answer i need to go

OpenStudy (festinger):

\(-47\frac{3}{4}\)

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (barinly):

thxs

OpenStudy (barinly):

@likeabossssssss can u help?

OpenStudy (lowkey.s):

HHAHHAA Barinly instead of brainly Nice name, mate! I encourage you to come here often there is a lot of awesome people and nice peps!

OpenStudy (lowkey.s):

Yes I can help!!

OpenStudy (barinly):

THXS :)

likeabossssssss (likeabossssssss):

@Festinger all ready aswered u @Barinly

OpenStudy (barinly):

its -191/4 but.. 191 is to big to be a numator

likeabossssssss (likeabossssssss):

ok im not sure what it is but ill tag some1 to help u ok

OpenStudy (barinly):

ok

OpenStudy (barinly):

@Festinger is it -47 3/4?

likeabossssssss (likeabossssssss):

@tshimp0629

likeabossssssss (likeabossssssss):

@Hayhayz can u help